Simplify 5[61-(9-99)]=
Answer by richwmiller(17219)

PEMDAS
parentheses exponents multiplication division addition subtraction
order of operations
first parentheses
start as far in as possible
(9-99)=-90
(61-(-90)
61+90=151
5*151=
you can finish it